OpenStudy (anonymous):

-8x^3 = 1 is x = 1/2 or -1/2?

OpenStudy (anonymous):

- 1/2

OpenStudy (anonymous):

Yeah, it's -1/2

OpenStudy (rogue):

Can a positive # cubed ever give you a negative #?

OpenStudy (anonymous):

to rogue, i dont think so... a positive will be positive even when cubed

hero (hero):

-8x^3 = 1 -1 = 8x^3 -1^3 = 8x^3 -1^3/8 = x^3 -1^3/2^3 = x^3 -1/2 = x

OpenStudy (anonymous):

i dont know what i did wrong in my question...

OpenStudy (anonymous):

the answer is positive..unless i did my derivative wrong y = -2x^4 so derivative is -8 x^3 is it correct?

OpenStudy (rogue):

This is a derivative question? Yes, the derivative of -2x^4 is -8x^3.

OpenStudy (anonymous):

and given the line is x- y +1=0

OpenStudy (anonymous):

so the slope for that is 1 right

OpenStudy (rogue):

What...? What is the question asking for?

OpenStudy (anonymous):

opps..nevermind.. it's asking for perpendicular slope! sorry !! thanks everyone :D

OpenStudy (rogue):

;] Good luck with the rest.